# ExPLoRA: Parameter-Efficient Extended Pre-Training to Adapt Vision Transformers under Domain Shifts

### v4

@@ Abstract
Parameter-efficient fine-tuning (PEFT) techniques such as low-rank adaptation (LoRA) can effectively adapt large pre-trained foundation models to downstream tasks using only a small fraction (0.1%-10%) of the original trainable weights. An under-explored question of PEFT is in extending the pre-training phase without supervised labels; that is, can we adapt a pre-trained foundation model to a new domain via efficient self-supervised pre-training on this domain? In this work, we introduce ExPLoRA, a highly effective technique to improve transfer learning of pre-trained vision transformers (ViTs) under domain shifts. Initializing a ViT with pre-trained weights on large, natural-image datasets such as from DinoV2 or MAE, ExPLoRA continues the unsupervised pre-training objective on a new domain, unfreezing 1-2 pre-trained ViT blocks and tuning all other layers with LoRA. We then fine-tune the resulting model only with LoRA on this new domain for supervised learning. Our experiments demonstrate state-of-the-art results on satellite imagery, even outperforming fully pre-training and fine-tuning ViTs. Using the DinoV2 training objective, we demonstrate up to 8% improvement in linear probing top-1 accuracy on downstream tasks while using <10% of the number of parameters that are used in prior fully-tuned state-of-the art approaches. Our ablation studies confirm the efficacy of our approach over other baselines such as PEFT. Code is available at: https://samar-khanna.github.io/ExPLoRA/
@@ 1 Introduction
In this work, we challenge this paradigm (fig. 1), asking whether pre-training from scratch on each new domain is strictly necessary, since doing so is expensive (in compute and time) and precludes knowledge transfer from natural images. Instead, we wish to more efficiently leverage the rich semantic information encoded in natural-image VFMs to adapt them to new domains. Our proposed solution addresses these concerns using parameter-efficient self-supervised learning methods for transfer learning.
@@
    Introducing ExPLoRA, a novel parameter-efficient method that extends unsupervised pre-training on target domains, achieving state-of-the-art supervised-learning performance using a fraction of the original ViT weights (section 5).
@@
    Conducting a comprehensive case study on satellite imagery, outperforming existing techniques on datasets like fMoW and showcasing improvements in linear probing accuracy. We also demonstrate generalization to multiple other domains within WILDS (section 6).
@@
    Demonstrating ExPLoRA’s efficacy via ablation studies and by analyzing improvements in local (positional) and global (class) information encoded in the patch representations output by each ViT block (section 7).
@@ 2 Related Work
Further comparisons with related work are in appendix A.
@@ 4 Problem Setup
Traditionally, the approach (fig. 1) has been to begin pre-training from scratch on data ${\mathcal{X}}_{T}$ from the new domain of interest, and then fine-tune for each dataset ${\mathcal{Y}}^{(\tau)}_{T}$, given as:
@@
The LoRA hypothesis is that the update $\Delta^{(\tau)}$ resides in a low-rank subspace when adapting pre-trained weights $W_{S}$ to fine-tuned weights $W^{(\tau)}_{T}$. This hypothesis holds well when pre-training and fine-tuning distributions are similar. However, when there is significant domain shift, such as between natural images and multi-spectral satellite data, the low-rank assumption often breaks down (see section 6.1.3).
@@
[…]
@@
[…]
@@ 5 Method
To learn $\Delta_{T}$, we propose ExPLoRA (i.e. Extended Pre-training with LoRA), a method that efficiently adapts a pre-trained ViT to a new target domain $T$, in algorithm 1.
@@
where $f_{\theta}(\cdot;W_{S})$ is a ViT feature encoder parameterized by $\theta$ with initialization $W_{S}$, $g_{\psi}$ is a learnable decoder parameterized by $\psi$ (e.g., MAE/DinoV2 decoder), ${\mathcal{C}}_{S}$ is the unsupervised loss that was used for $W_{S}$ (e.g., reconstruction loss), and $\Theta({\mathcal{U}},r)$ restricts the trainable parameter space to full-rank updates for blocks in ${\mathcal{U}}$ and LoRA rank-$r$ updates in ${\mathcal{L}}\setminus{\mathcal{U}}$. Similar to Goyal et al. (2023), we find that using the same unsupervised loss ${\mathcal{C}}_{S}$ as was used for $W_{S}$ is beneficial. As we show in section 6, ExPLoRA can even outperform full pre-training on new domains from scratch.
@@
For MAE, we initialize a ViT-L with $W_{S}$ from the MAE ViT-L encoder and use provided weights to initialize decoder $g_{\psi}$ (He et al., 2022). During extended pre-training, beyond the ExPLoRA recipe in algorithm 1, we apply LoRA with rank $r^{\prime}$ on the $Q,V$ matrices of each attention layer in $g_{\psi}$. The LoRA rank $r^{\prime}$ may differ from rank $r$ used in the ViT encoder $f_{\theta}$ (section B.4). All other decoder weights except layer-normalization are frozen, with no block fully unfrozen in $g_{\psi}$ to minimize trainable parameters.
For the multi-spectral ViT introduced by SatMAE we need to additionally unfreeze the positional encoding and patch embedding weights for each group of channels, as part of $\Theta({\mathcal{U}},r)$ in eq. 5. These cannot be initialized from $W_{S}$, as $W_{S}$ is trained on RGB inputs, whereas multi-spectral inputs can have more or different channels.
@@
After running ExPLoRA, we receive a new unsupervised model $W_{T}^{\ast}=W_{S}+\Delta_{T}$ for the target domain $T$ that functions as any other pre-trained ViT $W_{T}$. $g_{\psi}$ (e.g., the Dino/MAE decoder) is discarded as it is not part of the ViT encoder $f$. Only $\Delta_{T}$, consisting of 1-2 unfrozen ViT blocks, LoRA matrices, and layer-normalization weights, are stored for each $T$– all of which can be merged into the original ViT, thus preserving architecture. Like LoRA, ExPLoRA significantly reduces additional storage requirements compared to fully pre-training $W_{T}$.
@@
where the ViT $f_{\theta}$ is now initialized with $W_{T}^{\ast}$, $h^{(\tau)}_{\phi}(\cdot)$ is a task-specific ViT decoding head with parameters $\phi$, and ${\mathcal{C}}^{\prime}$ represents a supervised-learning loss function.
For any PEFT method, we allow $\theta\in\Theta(r)$, where $r$ restricts the trainable parameter space $\Theta$ of ${\mathcal{L}}$ (e.g., LoRA only on attention $Q,V$ matrices). For linear probing, $\Theta(0)=\emptyset$, so $(h^{(\tau)}_{\phi}\circ f_{\theta})({\mathbf{x}};W_{T}^{\ast})=\phi^{\top}f({\mathbf{x}};W_{T}^{\ast})$. For general fine-tuning, we optimize both $\theta\in\Theta$ and $\phi$ unrestricted, modifying $h^{(\tau)}$ as per the task (e.g., a decoder head for segmentation). With $\Delta^{(\tau)}$, eq. 4 gives us our final model weights $W^{(\tau)}_{T}$, which can be used for classification, segmentation, detection etc.
@@ 6 Experiments
Our experimental results consist of a case study on satellite imagery (section 6.1), with an ablation study in section 6.1.2. We evaluate on multiple downstream tasks in sections 6.1.3, 6.1.4 and 6.2. Additional experiments and ablations are provided in appendix B and training hyperparameter and compute configurations are mentioned in appendix C.
@@
Our results achieve a new SoTA top 1 accuracy of 79.3% ($\uparrow$1.5%) on the competitive fMoW-RGB benchmark, outperforming fully pre-trained and fine-tuned models while using 6% of the ViT encoder parameters and requiring only 100 GPU hours compared to 960+ hours for full pre-training. We also achieve a $\uparrow$8.2improvement in linear probing accuracy on the same dataset. Across other satellite datasets, ExPLoRA matches or exceeds fully-pretrained prior state-of-the-art methods while requiring 8x-10x less compute and 16x fewer trainable parameters and demonstrates competitive performance on WiLDS benchmark datasets as well.
@@ 6.1.1 RGB Satellite Images
We compare our results in table 1 against both prior fully pre-trained SoTA foundation models as well as PEFT techniques applied on ViTs pre-trained with MAE and/or DinoV2 weights. Our results demonstrate that D-ExPLoRA-$\tt{[L]}$-$r64$ is SoTA in terms of fMoW-RGB average accuracy at 79.28%. ExPLoRA outperforms techniques that require fully and/or continually pre-training ViTs on fMoW while using 6% of the original ViT encoder parameters and 8x less compute. Further experiments with MAE are in B.6.
@@
Using our strongest performing variant (i.e. ExPLoRA with DinoV2), we investigate linear-probing performance on fMoW-RGB compared with prior SoTA methods in table 2. Linear-probing represents freezing the backbone and then training a linear head on the features extracted from the frozen backbone, serving as a desirable metric of the quality of extracted embeddings. Our results demonstrate an improvement of over $\uparrow$8.2in top 1 average accuracy over prior SoTA methods, demonstrating that ExPLoRA learns robust unsupervised representations for its target domain without requiring expensive from-scratch pre-training. Importantly, ExPLoRA outperforms domain-specific prior SoTA solutions (rows 1-4), as well as DinoV2, which suggests successful transfer learning on the target domain by leveraging knowledge from pre-training on natural images.
@@ 6.1.2 Ablation study
We perform an ablation study (table 3) on linear-probing performance for fMoW-RGB to determine whether our proposed configuration performs optimally.
@@
[…]
Notably, full pre-training from scratch (row 2) achieves only 54.29% accuracy despite using 1200 GPU hours, suggesting that significantly more compute might be needed for a performant from-scratch DinoV2 checkpoint. In contrast, our standard configuration for ExPLoRA in row 15 achieves 77.48% accuracy with just 100 GPU hours (a 12$\times$ reduction in compute) while using only 18.7M parameters (6% of the full model). While rows 16-17 show further performance improvements with additional unfrozen blocks, row 15 represents our preferred trade-off between performance and efficiency for most applications. Further ablations on compute efficiency (B.2), data efficiency (B.3), MAE decoder rank (B.4), and ViT backbone size (B.5) are in appendix B.
@@ 6.1.3 Multi-Spectral Satellite Images
Table 4 highlights the challenge: direct fine-tuning of MAE weights on this domain results in a substantial 9% performance gap compared to SatMAE (rows 1 vs 2). LoRA tuning from MAE performs worse (row 3), and unfreezing four transformer blocks (row 7) fails to help. ExPLoRA bridges this gap effectively while requiring just 320 GPU hours and <10% trainable parameters for pre-training compared to 1150 hours for SatMAE. This demonstrates ExPLoRA’s ability to efficiently adapt to domains with substantial distribution shifts from natural images, preserving performance while dramatically reducing computational requirements.
@@ 6.1.4 Additional Satellite Datasets
Each input is a sequence of up to 3 fMoW-RGB (Christie et al., 2018) images of a location, distributed temporally, and paired with one of 62 classes. Since the inputs are now temporal sequences, we initialize the temporal MAE architecture from Cong et al. (2022) with MAE weights, and pre-train on ${\mathcal{X}}_{T}$ with ${\mathcal{U}}=\tt[L]$ and LoRA rank 32. ExPLoRA then outperforms temporal SatMAE for PEFT (table 5), demonstrating successful transfer learning at a fraction of the pre-training parameters and compute.
@@
This dataset contains high resolution satellite images, each paired with a segmentation mask for buildings (Van Etten et al., 2018). The training and test sets consist of 5000 and 1940 images, respectively. For ExPLoRA, we pre-train on the training set. However, many images in the dataset contain extensive blacked-out regions, indicating limits of the visible region. Considering this limitation and the small dataset size, it is not clear whether additional pre-training is effective. We find that, despite this, ExPLoRA remains on par with the LoRA-tuned DinoV2 model and remains competitive with the fully pre-trained and fully fine-tuned domain-specific models (table 6).
@@
The RESISC-45 (Cheng et al., 2017) benchmark dataset consists of 31,500 satellite images of varying resolution (0.2m-30m GSD), with 45 classes. The data is split into 25,200 training and 6,300 validation images, as per Reed et al. (2023). In table 6, our D-ExPLoRA pre-trained on only high-resolution fMoW-RGB images (last row of table 1) achieves SoTA results of 97.32% on multi-resolution RESISC-45 images, with just linear-probing. We demonstrate successful transfer learning from ExPLoRA pre-training, without requiring any additional modifications for scale-aware representation learning (Reed et al., 2023).

### v5

@@ Abstract
Parameter-efficient fine-tuning (PEFT) techniques such as low-rank adaptation (LoRA) can effectively adapt large pre-trained foundation models to downstream tasks using only a small fraction (0.1%-10%) of the original trainable weights. An under-explored question of PEFT is in extending the pre-training phase without supervised labels; that is, can we adapt a pre-trained foundation model to a new domain via efficient self-supervised pre-training on this domain? In this work, we introduce ExPLoRA, a highly effective technique to improve transfer learning of pre-trained vision transformers (ViTs) under domain shifts. Initializing a ViT with pre-trained weights on large, natural-image datasets such as from DinoV2 or MAE, ExPLoRA continues the unsupervised pre-training objective on a new domain, unfreezing 1-2 pre-trained ViT blocks and tuning all other layers with LoRA. We then fine-tune the resulting model only with LoRA on this new domain for supervised learning. Our experiments demonstrate state-of-the-art results on satellite imagery, even outperforming fully pre-training and fine-tuning ViTs. Using the DinoV2 training objective, we demonstrate up to 8% improvement in linear probing top-1 accuracy on downstream tasks while using <10% of the number of parameters that are used in prior fully-tuned state-of-the-art approaches. Our ablation studies confirm the efficacy of our approach over other baselines such as PEFT. Code is available at: https://samar-khanna.github.io/ExPLoRA/
@@ 1 Introduction
In this work, we challenge this paradigm (fig.˜1), asking whether pre-training from scratch on each new domain is strictly necessary, since doing so is expensive (in compute and time) and precludes knowledge transfer from natural images. Instead, we wish to more efficiently leverage the rich semantic information encoded in natural-image VFMs to adapt them to new domains. Our proposed solution addresses these concerns using parameter-efficient self-supervised learning methods for transfer learning.
@@
    Introducing ExPLoRA, a novel parameter-efficient method that extends unsupervised pre-training on target domains, achieving state-of-the-art supervised-learning performance using a fraction of the original ViT weights (section˜5).
@@
    Conducting a comprehensive case study on satellite imagery, outperforming existing techniques on datasets like fMoW and showcasing improvements in linear probing accuracy. We also demonstrate generalization to multiple other domains within WILDS (section˜6).
@@
    Demonstrating ExPLoRA’s efficacy via ablation studies and by analyzing improvements in local (positional) and global (class) information encoded in the patch representations output by each ViT block (section˜7).
@@ 2 Related Work
Further comparisons with related work are in appendix˜A.
@@ 4 Problem Setup
Traditionally, the approach (fig.˜1) has been to begin pre-training from scratch on data ${\mathcal{X}}_{T}$ from the new domain of interest, and then fine-tune for each dataset ${\mathcal{Y}}^{(\tau)}_{T}$, given as:
@@
The LoRA hypothesis is that the update $\Delta^{(\tau)}$ resides in a low-rank subspace when adapting pre-trained weights $W_{S}$ to fine-tuned weights $W^{(\tau)}_{T}$. This hypothesis holds well when pre-training and fine-tuning distributions are similar. However, when there is significant domain shift, such as between natural images and multi-spectral satellite data, the low-rank assumption often breaks down (see section˜6.1.3).
@@
1:Input: $W_{S}\coloneq$ pre-trained ViT with L layers ${\mathcal{L}}=\{\texttt{1},\dots,\texttt{L}\}$; ${\mathcal{X}}_{T}\coloneq$ unlabeled dataset
@@
3: Unfreeze all parameters of a subset of blocks $\mathcal{U}$$\subset{\mathcal{L}}$. (e.g., ${\mathcal{U}}=\{\texttt{L}\}$ or ${\mathcal{U}}=\{\texttt{1,L}\}$).
@@ 5 Method
To learn $\Delta_{T}$, we propose ExPLoRA (i.e. Extended Pre-training with LoRA), a method that efficiently adapts a pre-trained ViT to a new target domain $T$, in algorithm˜1.
@@
where $f_{\theta}(\cdot;W_{S})$ is a ViT feature encoder parameterized by $\theta$ with initialization $W_{S}$, $g_{\psi}$ is a learnable decoder parameterized by $\psi$ (e.g., MAE/DinoV2 decoder), ${\mathcal{C}}_{S}$ is the unsupervised loss that was used for $W_{S}$ (e.g., reconstruction loss), and $\Theta({\mathcal{U}},r)$ restricts the trainable parameter space to full-rank updates for blocks in $\mathcal{U}$} and LoRA rank-$r$ updates in ${\mathcal{L}}\setminus{\mathcal{U}}$. Similar to Goyal et al. (2023), we find that using the same unsupervised loss ${\mathcal{C}}_{S}$ as was used for $W_{S}$ is beneficial. As we show in section˜6, ExPLoRA can even outperform full pre-training on new domains from scratch.
@@
For MAE, we initialize a ViT-L with $W_{S}$ from the MAE ViT-L encoder and use provided weights to initialize decoder $g_{\psi}$ (He et al., 2022). During extended pre-training, beyond the ExPLoRA recipe in algorithm˜1, we apply LoRA with rank $r^{\prime}$ on the $Q,V$ matrices of each attention layer in $g_{\psi}$. The LoRA rank $r^{\prime}$ may differ from rank $r$ used in the ViT encoder $f_{\theta}$ (section˜B.4). All other decoder weights except layer-normalization are frozen, with no block fully unfrozen in $g_{\psi}$ to minimize trainable parameters.
For the multi-spectral ViT introduced by SatMAE we need to additionally unfreeze the positional encoding and patch embedding weights for each group of channels, as part of $\Theta({\mathcal{U}},r)$ in eq.˜5. These cannot be initialized from $W_{S}$, as $W_{S}$ is trained on RGB inputs, whereas multi-spectral inputs can have more or different channels.
@@
After running ExPLoRA, we receive a new unsupervised model $W_{T}^{\ast}=W_{S}+\Delta_{T}$ for the target domain $T$ that functions as any other pre-trained ViT $W_{T}$. $g_{\psi}$ (e.g., the Dino/MAE decoder) is discarded as it is not part of the ViT encoder $f$. Only $\Delta_{T}$, consisting of 1-2 unfrozen ViT blocks, LoRA matrices, and layer-normalization weights, are stored for each TT– all of which can be merged into the original ViT, thus preserving architecture. Like LoRA, ExPLoRA significantly reduces additional storage requirements compared to fully pre-training $W_{T}$.
@@
[…]
For any PEFT method, we allow $\theta\in\Theta(r)$, where $r$ restricts the trainable parameter space $\Theta$ of $\mathcal{L}$} (e.g., LoRA only on attention $Q,V$ matrices). For linear probing, $\Theta(0)=\emptyset$, so $(h^{(\tau)}_{\phi}\circ f_{\theta})({\mathbf{x}};W_{T}^{\ast})=\phi^{\top}f({\mathbf{x}};W_{T}^{\ast})$. For general fine-tuning, we optimize both $\theta\in\Theta$ and $\phi$ unrestricted, modifying $h^{(\tau)}$ as per the task (e.g., a decoder head for segmentation). With $\Delta^{(\tau)}$, eq.˜4 gives us our final model weights $W^{(\tau)}_{T}$, which can be used for classification, segmentation, detection etc.
@@ 6 Experiments
Our experimental results consist of a case study on satellite imagery (section˜6.1), with an ablation study in section˜6.1.2. We evaluate on multiple downstream tasks in sections˜6.1.3, 6.1.4 and 6.2. Additional experiments and ablations are provided in appendix˜B and training hyperparameter and compute configurations are mentioned in appendix˜C.
@@
Our results achieve a new SoTA top 1 accuracy of 79.3% ($\uparrow$1.5%) on the competitive fMoW-RGB benchmark, outperforming fully pre-trained and fine-tuned models while using 6% of the ViT encoder parameters and requiring only 100 GPU hours compared to 960+ hours for full pre-training. We also achieve a $\uparrow$8.2% improvement in linear probing accuracy on the same dataset. Across other satellite datasets, ExPLoRA matches or exceeds fully-pretrained prior state-of-the-art methods while requiring 8x-10x less compute and 16x fewer trainable parameters and demonstrates competitive performance on WiLDS benchmark datasets as well.
@@ 6.1.1 RGB Satellite Images
We compare our results in table˜1 against both prior fully pre-trained SoTA foundation models as well as PEFT techniques applied on ViTs pre-trained with MAE and/or DinoV2 weights. Our results demonstrate that D-ExPLoRA-$\tt{[L]}$-$r64$ is SoTA in terms of fMoW-RGB average accuracy at 79.28%. ExPLoRA outperforms techniques that require fully and/or continually pre-training ViTs on fMoW while using 6% of the original ViT encoder parameters and 8x less compute. Further experiments with MAE are in B.6.
@@
Using our strongest performing variant (i.e. ExPLoRA with DinoV2), we investigate linear-probing performance on fMoW-RGB compared with prior SoTA methods in table˜2. Linear-probing represents freezing the backbone and then training a linear head on the features extracted from the frozen backbone, serving as a desirable metric of the quality of extracted embeddings. Our results demonstrate an improvement of over $\uparrow$8.2% in top 1 average accuracy over prior SoTA methods, demonstrating that ExPLoRA learns robust unsupervised representations for its target domain without requiring expensive from-scratch pre-training. Importantly, ExPLoRA outperforms domain-specific prior SoTA solutions (rows 1-4), as well as DinoV2, which suggests successful transfer learning on the target domain by leveraging knowledge from pre-training on natural images.
@@ 6.1.2 Ablation study
We perform an ablation study (table˜3) on linear-probing performance for fMoW-RGB to determine whether our proposed configuration performs optimally.
A natural question is whether the improvement in performance stems primarily from unfreezing blocks, or from LoRA-tuning the rest of the ViT. We investigate this by unfreezing blocks $\{\texttt{L,L-1}\}$ (with no LoRA) in row 4, and comparing that with ExPLoRA-L-$r8$ in row 13. As seen, unfreezing an extra block consumes almost double the number of parameters, but fails to yield the same improvement in performance $\downarrow 0.34\%$. Thus, simply increasing the number of unfrozen blocks will likely improve performance, but will not do so as effectively as ExPLoRA, and will also significantly and sharply decrease the parameter-efficiency.
The choice of $\mathcal{U}$} matters as well. As seen in rows 8-10, and 15, for the DinoV2 objective, ${\mathcal{U}}=\{\texttt{1}\}$ or ${\mathcal{U}}=\{\texttt{9}\}$ are not as effective as ${\mathcal{U}}=\{\texttt{L-1}\}$ or ${\mathcal{U}}=\{\texttt{L}\}$, ceteris paribus. To understand this result further, see section˜7. We also notice a slight drop in accuracy from leaving the normalization layers across the ViT frozen, seen in row 12.
Lastly, we investigate the impact of LoRA rank and increasing the set of unfrozen blocks $\mathcal{U}$} on ExPLoRA, effectively varying the parameter budget $\Theta({\mathcal{U}},r)$. As seen in table˜3, changing the rank from 8 to 32 yields a modest improvement ($\uparrow 0.09\%$, rows 13 vs 14), but increasing from 32 to 64 brings about a much larger gain ($\uparrow 1.08\%$, rows 14 vs 15), with an equivalent 3M parameter increase. Meanwhile, expanding the set of unfrozen blocks from ${\mathcal{U}}=\{\texttt{L}\}$ to ${\mathcal{U}}=\{\texttt{L-1, L}\}$ (rows 15 vs 16) and further to ${\mathcal{U}}=\{\texttt{1, L-1, L}\}$ (rows 16 vs 17) yield consistent improvements of $\uparrow 0.28\%$.
@@ 6.1.3 Multi-Spectral Satellite Images
Table˜4 highlights the challenge: direct fine-tuning of MAE weights on this domain results in a substantial 9% performance gap compared to SatMAE (rows 1 vs 2). LoRA tuning from MAE performs worse (row 3), and unfreezing four transformer blocks (row 7) fails to help. ExPLoRA bridges this gap effectively while requiring just 320 GPU hours and <10% trainable parameters for pre-training compared to 1150 hours for SatMAE. This demonstrates ExPLoRA’s ability to efficiently adapt to domains with substantial distribution shifts from natural images, preserving performance while dramatically reducing computational requirements.
@@ 6.1.4 Additional Satellite Datasets
Each input is a sequence of up to 3 fMoW-RGB (Christie et al., 2018) images of a location, distributed temporally, and paired with one of 62 classes. Since the inputs are now temporal sequences, we initialize the temporal MAE architecture from Cong et al. (2022) with MAE weights, and pre-train on ${\mathcal{X}}_{T}$ with ${\mathcal{U}}=\tt[L]$ and LoRA rank 32. ExPLoRA then outperforms temporal SatMAE for PEFT (table˜5), demonstrating successful transfer learning at a fraction of the pre-training parameters and compute.
@@
This dataset contains high resolution satellite images, each paired with a segmentation mask for buildings (Van Etten et al., 2018). The training and test sets consist of 5000 and 1940 images, respectively. For ExPLoRA, we pre-train on the training set. However, many images in the dataset contain extensive blacked-out regions, indicating limits of the visible region. Considering this limitation and the small dataset size, it is not clear whether additional pre-training is effective. We find that, despite this, ExPLoRA remains on par with the LoRA-tuned DinoV2 model and remains competitive with the fully pre-trained and fully fine-tuned domain-specific models (table˜6).
@@
The RESISC-45 (Cheng et al., 2017) benchmark dataset consists of 31,500 satellite images of varying resolution (0.2m-30m GSD), with 45 classes. The data is split into 25,200 training and 6,300 validation images, as per Reed et al. (2023). In table˜6, our D-ExPLoRA pre-trained on only high-resolution fMoW-RGB images (last row of table˜1) achieves SoTA results of 97.32% on multi-resolution RESISC-45 images, with just linear-probing. We demonstrate successful transfer learning from ExPLoRA pre-training, without requiring any additional modifications for scale-aware representation learning (Reed et al., 2023).